the measure of the second angle of a triangle is twice as great as that of the first angle. If the measure of the thrid angle is 20 degrees greater than that of the second angle, what is measure of the third angle
Answer by ptaylor(2198)


Let x=the measure of the first angle
Then 2x=the measure of the second angle
And (2x+20)=the measure of the third angle
Now we know that all three angles must add up to 180 degrees. So:
x+2x+(2x+20)=180 get rid of parens
x+2x+2x+20=180 subtract 20 from both sides
x+2x+2x+20-20=180-20 collect like terms
5x=160 divide both sides by 5
x=32 degrees--------------------first angle
2x+20=2(32)+20=64+20= 84 degrees third angle-----------------------ans
2x=2(32)=64 degrees ------------second angle
CK
84+64+32=180
180=180
also 64 equals two times 32 and 84 equals 2 times 32 +20

Hope this helps----ptaylor